Find the center-radius form of the circle whose equation is
x2 + 6x + y^2 - 2y + 6 = 0.

Answer by mananth(16949)

x2 + 6x + y^2 - 2y + 6 = 0.
..
(x^2+6x+9)-9 +(y^2-2y+1)-1 +6=0
..
(x+3)^2 +(y-1)^2 -9-1+6=0
(x+3)^2 +(y-1)^2 = 4
...
Center (-3,1)
radius = sqrt4 =2
...
m.ananth@hotmail.ca